OpenStudy (anonymous):

If (5 + 1)(6 ÷ 3)(8 − 5) = (3 + 3)x, then x = A. 12 B. 3 C. 4 D. 6

OpenStudy (anonymous):

I keep getting a decimal and I don't know what im doing wrong if you solve this could you please show a little work so I know what im doing wrong thanks!

OpenStudy (anonymous):

(5+1)(6/3)(8-5)=(3+3)x (6)(2)(3)=6x 36=6x 36/6=6/6x 6=x x=6

OpenStudy (anonymous):

Okay thanks!!! :)

OpenStudy (anonymous):

just use PEMDAS